OpenStudy (anonymous):

what is an equation of the line that passes through the points (1,3) and (8,5)

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

y = 2/7x + 2

OpenStudy (anonymous):

or in point slope y - 3 = 2/7 (x - 1)

OpenStudy (amorfide):

first you need to work out the slope there formula is \[m=\frac{ y_{2}-y_{1} }{ x_{2}-x_{1} }\] where m is the slope (x1,y1)=(1,3) (x2,y2)=(8,5) now we must work out the equation of the line the formula is \[y-y_{1}=m(x-x_{1})\]